OpenStudy (harrisonshark):

Simplify the following expression. 1 over z^-3

OpenStudy (harrisonshark):

help fast please

OpenStudy (anonymous):

if you recall my answer for your previous question which was \[6^{-3}\]the result you got was \[\frac{ 1 }{ 6^3 }\]i mentioned that this happened because in order to get rid of a negative exponent, you need to put the number being raised to the bottom of a fraction, the same concept applies so you'll get that \[\frac{ 1 }{ z^{-3} }=\frac{ 1 }{ \frac{ 1 }{ z^3 } }=z^3\]because when dividing by a fraction, you multiply the numerator by the reciprocal (flipped) denominator

OpenStudy (harrisonshark):

thank you so much

OpenStudy (harrisonshark):

:*

OpenStudy (harrisonshark):

pellet i forgot im on my brothers account hahahahahah